kotok123456:

The circumference of a circle is 19.468 meters. Find the radius of the circle. 3.1 m 6.2 m

jhonyy9:

formula for circumference of a circle is C = 2*pi*r => r = C/(2pi) given C = 19.468 m r = 19468/(2*3,14) = make the calcule hope this will help you